OpenStudy (anonymous):

Which set of numbers could represent the lengths of the sides of a right triangle? Select one: a. 5, 12, 13 b. 12, 13, 15 c. 10, 12, 14 d. 10, 12, 22

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

Do you know the theorem that applies to right triangles?

OpenStudy (anonymous):

No

myininaya (myininaya):

The hint to apply here is the Pythagorean Theorem.

myininaya (myininaya):

a^2+b^2=c^2 where c is the bigger of the 3 numbers

OpenStudy (anonymous):

ohh okay I can do that

myininaya (myininaya):

so have you decided which of the 4 sets above satisfy a^2+b^2=c^2?

myininaya (myininaya):

which of the following are true: \[5^2+12^2=13^2\\ 12^2+13^2=15^2\\ 10^2+12^2=14^2\\ 10^2+12^2=22^2\]

OpenStudy (anonymous):

@myininaya the first ones true : 5^2 +12^2=13^2

myininaya (myininaya):

cool stuff